OpenStudy (anonymous):

help me omg!! which expression gives the solution of: -5+2x(2)=-6x? see replies!!!

OpenStudy (anonymous):

a. \[2\pm \sqrt{4-(4)(6)(-5)} \over 12\]

OpenStudy (anonymous):

b. \[-5\sqrt{25-(4)(2)(6)} \over -10\]

OpenStudy (anonymous):

c. \[-6\pm \sqrt{36-(4)(2)(-5)} \over 4\]

OpenStudy (anonymous):

d.\[6\pm \sqrt{36-(4)(2)(-5)}\]

OpenStudy (anonymous):

c.

OpenStudy (anonymous):

can you explain why please? i have to show some work.

OpenStudy (anonymous):

-5+2x^2+6x=0 \[\sqrt{-b \pm b^2 -4ac} / 2a\]

OpenStudy (anonymous):

b=6 a=2 c=-5

OpenStudy (anonymous):

\[2x ^{2} + 6x - 5 = 0\] By the quadratic formula, \[x= -6 \pm \sqrt{6^{2} - (4)(2)(-5)} \div 4\] The quadratic formula is \[x = -b \pm \sqrt{b ^{2} - 4ac} \div 2a \] for \[ax ^{2} + bx + c =0\]

OpenStudy (anonymous):

thank you so much. :*

OpenStudy (anonymous):

Welcome!